# Constraint Learning for Parametric Point Cloud

### v3

@@ 3.5 Constraint Feature Learning
Stage 2 of CstNet is designed to leverage constraints to deepen the model’s understanding of CAD shapes. The designed architecture extracts features from Primitive Type, Main Axis Direction, and Adjacency separately, attention between features is then applied to update constraint feature. The structure and module details are illustrated in Fig. 6.
The Triple MLP takes the point cloud, Primitive Type, Main Axis Direction, and Adjacency as inputs. In this module, the point cloud is concatenated with the Primitive Type, Main Axis Direction, and Adjacency separately before being fed into three individual MLPs. The outputs are the Primitive Feature, Axis Feature, and Adjacency Feature, respectively.
In the C-MLP layer, point cloud is concatenated with Primitive Type, Main Axis Direction, Adjacency, and then feeds into MLP, with the output being the Parametric Feature.
@@
The Fea Attention layer also takes the Primitive Feature, Axis Feature, Adjacency Feature, and Parametric Feature as input, attention between features is employed to update the constraint feature efficiently. This design considers that different points may focus on different features, for example, points near edge may focus more on the Adjacency Feature, while others may pay more attention to the Primitive or Axis Feature. The Fea Attention is developed by vector attention as formulated in Eq. 2.
$\boldsymbol{f}^{\prime}_{cst}=\sum_{\boldsymbol{f}_{j}\in\mathcal{F}}\rho\Bigl{(}mlp\bigl{(}\mathrm{Q}(\boldsymbol{f}_{cst})-\mathrm{K}(\boldsymbol{f}_{j})\bigr{)}\Bigr{)}\odot\mathrm{V}(\boldsymbol{f}_{cst}).$ (2)
In which =$\mathcal{F}=\{\boldsymbol{f}_{pmt},\boldsymbol{f}_{axs},\boldsymbol{f}_{adj},\boldsymbol{f}_{cst}\}$, $\boldsymbol{f}_{pmt}$ is Primitive Feature, $\boldsymbol{f}_{axs}$ is Axis Feature, $\boldsymbol{f}_{adj}$ is Adjacency Feature, and $\boldsymbol{f}_{cst}$ is Constraint Feature. $\Delta$ is defined as one-hot encoding of $\boldsymbol{f}_{j}$ type.
For downstream tasks, if classification is required, the final constraint feature can be fed into MLP and softmax. For segmentation tasks, upsampling and skip link concatenation similar to those in PointNet++ [34] can be applied.
@@ 4.1 Experimental Settings
In general, we train CstNet using the Negative Log Likelihood Loss (NLLLoss) and optimized by Adam optimization for 200 epochs, with a weight decay of 0.0001 and an initial learning rate of 0.0001. We use StepLR learning rate scheduler, with step size of 20 and gamma of 0.7. The batch size is 16. All experiments are performed on GeForce RTX 4090 GPU.
Parametric 20000 dataset: our Parametric 20000 dataset contains 19,739 multi-modal instances categorized into 75 classes. Each instance includes B-Rep data, mesh, and point cloud. All B-Rep data are collected from the following three sources. More details in the supplementary.
Sources 1 (36.05 %): Downloaded and purified from TraceParts [16]. The CAD shapes on TraceParts are collected from manufacturing companies.
Sources 2 (24.12 %): Provided by a CAD company [^1] https://www.newdimchina.com/ that collaborates with our team. These shapes are all mechanical components.
Sources 3 (39.83 %): Instantiated from parametric templates designed by our team. These templates are designed based on common mechanical structures.
Our Parametric 20000 dataset encompasses various common formats of CAD shapes, making it suitable for a wide range of tasks. The careful selection and purification process undertaken by our team further ensures its high quality.
@@ 4.2 Constraint Acquisition
The extraction of constraints from B-Rep data is accomplished by CST-BRep module, with results visualized in the "GT" column of Fig. 8. By referring to the model row, this module performs well in extracting three components of constraint.
Extracting constraints from point cloud is achieved by CST-PCD module. The training set is sourced from 25 trunks of ABC [18], with each trunk containing 10,000 B-Rep files. We utilized CST-BRep module to extract constraints as Ground Truth for model training. After the training finished, the prediction accuracy of Primitive Type reached 94.23 %, for Adjacency, that is 94.62 %, and the MSE loss for Main Axis Direction is 0.0559.
In subsequent classification and segmentation tasks, the above pre-trained CST-PCD module is utilized to predict constraints with its weights frozen and no further updates. To ensure that only point cloud is used as input across all methods in comparison, even when B-Rep data is available, the CST-BRep module will not be activated to obtain constraints.
After CST-PCD module was pre-trained, we evaluated it on a new trunk of ABC. The visualization of constraint prediction is shown in the "Pred." column of Fig. 8. After traversing this trunk, the prediction accuracy of Primitive Type reached 91.19 %, for Adjacency, that is 90.24 %, the MSE loss of Main Axis Direction is 0.0998, above demonstrating the strong capability for constraint prediction.
Figure 8: Constraint prediction on ABC. MAD: Main Axis Direction. PT: Primitive Type. ADJ: Adjacency (whether point is near edge). Main Axis Direction has been processed for visualization, details in supplementary.
Additionally, we evaluated the pre-trained CST-PCD module on the CAD shape dataset MCB [16] to predict constraints. For the MCB consists of mesh files, it was not possible to extract Ground Truth constraints for comparison. Based on the results visualized in Fig. 9, CST-PCD demonstrated a high level of constraint prediction accuracy even on an unseen dataset.
Figure 9: Constraint prediction on MCB.
@@ 4.3 Classification
The classification results on the Parametric 20000 dataset are presented in Tab. 1, where our CstNet outperformed other methods across both metrics. Additionally, the instance accuracy of all methods is below 90 %, with a notable performance gap between them, indicating that this dataset poses challenges and effectively differentiates between methods.
Table 1: Classification results on Parametric 20000. Acc: accuracy over instance %, Acc*: accuracy over class %, F1: F1-score, mAP: mean average precision %.
@@
PointCNN [24] 61.92 55.29 53.97 40.99
PointNet [33] 75.61 74.08 74.23 82.15
PointNet++ [34] 82.89 86.80 82.31 88.05
PointConv [47] 80.53 82.10 81.86 86.54
DGCNN [42] 85.99 88.20 87.20 89.20
3DGCN [26] 86.19 86.31 85.37 87.64
Ours 89.93 91.04 90.32 92.32
We conducted evaluations on the CAD shape dataset MCB [16], with results presented in Tab. 2. Our CstNet achieved the highest scores across all metrics on MCB-A, but performed less effectively on MCB-B. This discrepancy may stem from the composition of MCB: MCB-A includes shapes from TraceParts (69.27 %), 3D Warehouse (21.70 %), and GrabCAD (9.03 %), while MCB-B is limited to shapes from 3D Warehouse and GrabCAD. TraceParts provides data from manufacturing companies, whereas 3D Warehouse and GrabCAD contain shapes for general usage [16]. As the CST-PCD module in CstNet is pre-trained on ABC, which primarily comprises mechanical components, this pre-training may lead to poor constraint prediction performance on MCB-B, thereby influencing the classification results.
Table 2: Classification results on MCB. A / B column: results of MCB A / B subset.
Method Acc.instance % Acc.class % F1-score mAP %
A B A B A B A B
PointCNN [24] 93.89 93.67 81.85 86.80 83.86 88.63 90.13 93.86
PointNet [33] 86.78 80.36 67.70 65.30 86.55 79.89 74.08 71.66
PointNet++ [34] 87.45 93.91 73.68 87.97 88.32 73.45 91.33 91.37
SpiderCNN [53] 93.59 89.31 79.70 79.29 81.30 80.72 86.64 82.47
PointConv [47] 93.25 90.64 80.24 80.17 71.31 76.07 82.19 86.09
DGCNN [42] 92.54 90.75 74.47 78.17 76.12 79.69 74.27 79.73
3DGCN [26] 93.71 92.13 78.71 84.31 84.59 79.24 84.35 84.35
Ours 96.87 92.61 89.21 85.94 89.85 85.68 93.17 91.25
4.4 Segmentation
We conducted experiments on the mechanical segmentation dataset 360 Gallery [19], which contains point clouds with segmentation labels based on CAD modeling operations. However, this labeling approach presents disadvantages for point cloud segmentation. For example, a cube in this dataset has four faces labeled as "ExtrudeSide" and two faces labeled as "ExtrudeEnd", which poses an unrealistic task for point cloud networks. Additionally, the distribution of points across different labels is highly imbalanced. For instance, points labeled as "ExtrudeSide" account for 45.94 %, while those labeled as "RevolveEnd" make up only 0.12 %. Despite these limitations, to our knowledge, 360 Gallery is currently the only publicly available point cloud dataset for CAD shape segmentation tasks.
The segmentation results are presented in Tab. 3, where our method outperforms the state-of-the-art methods on most metrics. The results also highlight the limitations of 360 Gallery in terms of point cloud segmentation. Apart from "ExtrudeSide" and "ExtrudeEnd", none of the models were able to differentiate the other segmentation categories except DGCNN managed to predict "RevolveSide" with a low mIoU.
Table 3: Part Segmentation Results on 360 Gallery. PN: PointNet. PN2: PointNet++. DGN: DGCNN. 3DN: 3DGCN. OA: overall accuracy %. mIoU: instance mIoU %. mIoU*: class mIoU %. ESide: Extrude Side mIoU %, EEnd: Extrude End mIoU %. Other segmentation categories’ mIoU are all 0 except RevolveSide mIoU of DGCNN is 9.60 %, more datails in supplementary.
Method OA mIoU mIoU* ESide EEnd
PN [33] 42.13 18.89 7.04 19.83 36.47
PN2 [34] 58.18 26.98 11.27 49.80 40.36
DGN [42] 64.02 35.54 14.36 54.48 50.82
3DN [26] 60.58 33.03 12.23 50.67 47.20
Ours 70.76 46.65 20.51 61.29 60.65
A portion of the segmentation results on 360 Gallery test set are shown in Tab. 4, where our CstNet shows segmentation results closest to the Ground Truth. PointNet’s segmentation differs from other methods, for instance, in the third and fourth rows, PointNet mistakenly segments a single primitive into two different categories. This discrepancy likely arises from the model’s inability to effectively capture local features.
Table 4: Visualization of part segmentation on 360 Gallery. GT: ground truth. PN: PointNet [33]. PN2: PointNet++ [34]. DGN: DGCNN [42]. 3DN: 3DGCN [26].
GT PN PN2 DGN 3DN Ours
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ 4.5 Ablation Studies
Our model leverages all three components of constraint: Primitive Type, Main Axis Direction, and Adjacency. Additionally, we utilized the SurfaceKNN to obtain more suitable neighbors for each point. To validate the effectiveness of our design, we conducted a series of experiments on Parametric 20000 dataset.
For each component of constraint, we evaluated their importance by disabling one at a time. This operation is performed by setting it as a zero tensor before feeding into the model. The results are presented in Tab. 5, which show that all three components contribute to improving the performance of our CstNet.
To assess the effectiveness of SurfaceKNN, we compared it with KNN. From Tab. 5, it can be concluded that in most cases, our SurfaceKNN outperforms KNN. However, when the Primitive Type is disabled, the results are similar between the two methods, indicating that SurfaceKNN primarily enhances the effectiveness of the Primitive Feature.
Table 5: Effects on Parametric 20000 dateset with varying neighbor search and constraints. Row 3 - 6: SurfaceKNN, Row 7 - 10: KNN.
@@ 5 Conclusions
In this paper, we studied the CAD shapes and introduced the constraints expression which is suitable for deep learning methods. Afterward, we proposed the CstNet for extracting and learning constraint features. Additionally, we built up a multimodal CAD shape classification dataset Parametric 20000. Experimental results on both the public datasets and proposed dataset demonstrate that our CstNet achieves state-of-the-art performance across multiple metrics. Comprehensive ablation studies further validate that each component of constraint and SurfaceKNN contributes to improve the performance of CstNet.
CAD shapes possess numerous attributes. In this paper, we analyze the constraints from a functional perspective. In future work, we plan to investigate additional aspects such as part design and manufacturing processes, to achieve a deeper understanding of CAD shapes.
@@ A Overview
In Sec. B, we present the details of network architecture and data preparation in validation experiments. In Sec. D, we show more details of our constraint representation and constraint learning methods. In Sec. C, we provide the data construction and static of Parametric 20000 dataset. Finally, Sec. E offers data process for visualization and more results of experiments.
B Validation Experiment
B.1 Constrain Expression
@@ B.3 Model Design
C Parametric 20000 Dataset
C.1 Dataset Construction
This section provide details of purification process and data acquisition of Parametric 20000 dataset.
@@ C.1 Dataset Construction
Point cloud: The point cloud data in Parametric 20000 dataset containing about 2,200 points, point constraints is also included, each point is expressed as (Coordinate, Primitive Type, Main Axis Direction, Adjacency), the meshes and point clouds are generated from B-Rep data.
C.2 Dataset Statics
This section presents the shapes count and proportion statistics for each category in the Parametric 20000 dataset. Additionally, shapes from each category are showcased.
Parametric 20000 dataset is divided into 80 % for training and 20 % for testing.

### v5

@@ 3.4 Constraint-Aware Feature Learning
The Stage 2 of CstNet is designed to leverage constraints effectively, which extracts features from MAD, Adj, PT separately, and attention between features is then applied to update constraint feature, with details in Fig. 7.
@@
$\boldsymbol{f}^{\prime}_{\rm{cst}}=\sum_{\boldsymbol{f}_{j}\in\mathcal{F}}\rho\Bigl{(}\rm{MLP}\bigl{(}\mathrm{Q}(\boldsymbol{f}_{\rm{cst}})-\mathrm{K}(\boldsymbol{f}_{\mathit{j}})\bigr{)}\Bigr{)}\odot\mathrm{V}(\boldsymbol{f}_{\rm{cst}}).$ (2)
In which =$\mathcal{F}=\{\boldsymbol{f}_{\rm{mad}},\boldsymbol{f}_{\rm{adj}},\boldsymbol{f}_{\rm{pt}},\boldsymbol{f}_{\rm{cst}}\}$, corresponding to Axis Feature, Adjacency Feature, Primitive Feature, and Constraint Feature. $\Delta$ is the one-hot encoding of $\boldsymbol{f}_{j}$ type.
For downstream classification tasks, the constraint features obtained in each module will be concatenated by residual connection and feed into maxpoling, MLPs and softmax. Constraint features obtained by different modules may differ in features’ number and channel, based on the coordinates of the final constraint features, we identify the correspondence from different modules for further processing.
@@ 4.1 Experimental Settings
We train CstNet using the Negative Log Likelihood Loss and optimized by Adam optimization for 200 epochs, with a weight decay of 0.0001 and an initial learning rate of 0.0001. We use StepLR learning rate scheduler, with step size of 20 and gamma of 0.7. The batch size is 16. All experiments are performed on GeForce RTX 4090 GPU.
Param20K dataset Param20K contains 19,739 multi-modal instances categorized into 75 classes. Each instance includes B-Rep data, mesh, and point cloud. All B-Rep data are collected from the following three sources. More details in the supplementary. Sources 1 (36.05 %): Downloaded and purified from TraceParts [18]. The CAD shapes on TraceParts are collected from manufacturing companies. Sources 2 (24.12 %): Provided by a CAD company [^1] https://www.newdimchina.com/ that collaborates with our team. These shapes are all mechanical components. Sources 3 (39.83 %): Instantiated from parametric templates designed by our team. These templates are designed based on common mechanical structures.
Our Param20K dataset encompasses various common formats of CAD shapes, making it suitable for a wide range of tasks. The careful selection and purification process undertaken by our team further ensures its high quality.
@@ 4.2 Constraint Acquisition
The extraction of constraints from B-Rep data is accomplished by the CstBRep module, with results visualized in the Ground Truth column of Fig. 8. By referring to the model row, CstBRep performs well in extracting MAD, Adj, and PT from B-Rep data. Predicting MAD, Adj, and PT from point cloud is achieved by the CstPnt module. The training set is sourced from 25 trunks of ABC [20], with each trunk containing 10,000 B-Rep files. We utilized CstBRep module to extract constraints as Ground Truth for model training. After the training finished, the MSE loss of MAD is 0.0559, the prediction accuracy of Adj and PT reached 94.62% and 94.23%, respectively. In subsequent experiments, the above pre-trained CstPnt module is utilized to predict MAD, Adj, and PT with its weights frozen and no further updates. To ensure that only point cloud is used as input across all methods when comparing, even when B-Rep data is available, the CstBRep module will not be activated to obtain MAD, Adj, and PT.
After CstPnt was pre-trained, we evaluated it on a ABC new trunk and Param20K dataset. The ABC new trunk was treated as the test set, while the Param20K was considered as an unseen CAD shapes dataset, more results on other public available CAD shapes datasets in supplementary. Since there is no existing model that directly predicts MAD, Adj, and PT, we modified the published CAD-specific methods [58, 41] by adding MLPs as prediction head to enable these predictions. The experiment results are shown in Tab. 1, from which our CstPnt achieves the highest accuracy in predicting MAD, Adj, and PT. Furthermore, from ABC to Param20K, all methods’ performance decreased, but our method dropped least accuracy. ParSeNet and HPNet may be negatively affected by global features, which impaired their performance on unseen datasets. Fig. 8 presents the visualization of constraint prediction results, which shown that our method is closest to the ground truth. Furthermore, from ABC to Paeam20K, our method demonstrates excellent generalization performance on unseen datasets. In contrast, while ParSeNet and HPNet exhibit acceptable prediction results on the ABC new trunk, they performed unsatisfactory on the unseen Param20K dataset.
Table 1: Constraint prediction on ABC (new trunk) and Param20K. MAD: MSE loss of Main Axis Direction. Adj, PT: prediction accuracy of Adjacency and Primitive Type.
[…]
ABC ParSeNet [41] 0.1356 84.89 85.22
HPNet [58] 0.1576 82.63 79.15
Ours 0.0998 90.24 91.19
Param20K ParSeNet [41] 0.2247 81.42 59.75
HPNet [58] 0.2570 78.60 57.66
Ours 0.1390 87.95 86.52
Figure 8: Constraint prediction visualization on ABC and Param20K. MAD is processed for visualization, details in supplementary.
@@ 4.3 Classification
To validate the ability of our CstNet in identifying CAD shapes that are visual similar but different in constraints, we conducted experiments on the Prism-Cuboid datasets presented in the Problem Statement section Sec. 3.1. Results shown in Fig. 9, in which our CstNet achieved best instance accuracy in distinguish 50°50°50  ° to 87°87°87  ° prisms and cuboids, which demonstrate the strong ability of CstNet in recognizing shapes with similar appearance but constraints different. For the binary classification of 89°89°89  ° prisms and cuboids, all methods achieved approximately 50% instance accuracy, this is attributed to the differences between them are so subtle that no one is capable to distinguishing them.
Figure 9: Classification results on Prism-Cuboid dataset.
The classification of prisms and cuboids is predominantly influenced by constraints, while the comprehension of common CAD shapes necessitates the integration of both constraint and geometric features. CstNet achieves this by concatenating XYZ with constraint in Stage 2, thereby integrating geometric features into constraint features. To evaluate CstNet’s performance on common CAD shapes, we conducted experiments on Param20K dataset (more results on public CAD shapes dataset in Supplementary), results are presented in Tab. 2, where our CstNet outperformed other methods across both metrics, validated the effectiveness of utilizing constraint and geometric features.
Table 2: Classification results on Param20K dataset. Acc: accuracy over instance %, Acc*: accuracy over class %, F1: F1-score, mAP: mean average precision %.
@@
PointCNN [27] 80.25 76.03 74.64 83.38
PointNet [36] 81.30 83.21 82.06 85.18
PointNet++ [37] 83.70 86.37 85.30 87.94
PointConv [51] 82.30 85.62 84.91 87.92
DGCNN [46] 85.40 87.28 86.43 89.17
3DGCN [29] 86.42 88.25 87.63 89.81
Ours 89.94 91.06 90.34 92.72
To access the robustness of CstNet, we evaluated it under rotated point clouds. Since point clouds are pre-processed by aligning their centroids to the origin and scaling their ranges into a unit cube, the performance of all methods is unaffected by point cloud shifts and scaling. Comparison results depicted in Fig. 10, from which our CstNet demonstrates strongest robustness to point cloud rotation. The dashed line in Fig. 10 show that the performance of CstPnt module remains nearly unaffected by rotation, owing to it learned various local features from ABC dataset.
Figure 10: Classification results on rotated Param20K. The training set remained unchanged and the test set was rotated by a certain angle along the +Z direction. Solid line: left y axis, dashed line: right y axis. Acc.MAD = 1 - MAD.MSELoss.
@@ 4.4 Ablation Studies
Our CstNet leverages all three components of constraint: MAD, Adj, and PT. Additionally, we utilized the SurfaceKNN to obtain more efficient neighbors for each point. To validate the effectiveness of our design, we conducted a series of ablation studies on Param20K dataset.
For each component of constraint, we evaluated their importance by disabling one or all at a time. This operation is performed by setting the disabled components as a zero tensor before feeding into CstNet. Classification results are presented in Tab. 3, which show that all three components contribute to improving the performance of our CstNet. To assess the effectiveness of SurfaceKNN, we compared it with KNN. From Tab. 3, it can be concluded that in most cases, our SurfaceKNN outperforms KNN. However, when the Primitive Type is disabled, the results are similar between the two methods, indicating that SurfaceKNN primarily enhances the effectiveness of the Primitive Feature.
Table 3: Effects on varying neighbor search and constraint components. Upper half: SurfaceKNN, lower half: KNN.
@@
83.19 82.54 81.46 86.34
To validate the effectiveness of Stage 2 backbone, we conducted a comprehensive evaluation by feeding (x, y, z, MAD, Adj, PT) into multiple classification backbone for comparison. The results are presented in Tab. 4, from which our method demonstrates best performance under both predicted and label constraints, indicating that our design achieves the best utilization of constraints. Furthermore, although the accuracy using predicted constraints is lower than that of using label constraints, it still shows a significant improvement compared to the baselines without constraints. This highlights the effectiveness of the predicted constraints in improving model performance.
Table 4: Effects on Stage2 Backbone. Upper half: predicted constraints from CstPnt, lower: label constraints.
Backbone-S2 Acc Acc* F1 mAP
PointNet [36] 83.83 83.64 83.20 87.27
PointNet++ [37] 85.52 88.11 87.18 90.79
DGCNN [46] 86.81 88.70 87.70 91.35
3DGCN [29] 87.46 89.59 87.97 91.92
Ours 89.94 91.06 90.34 92.72
PointNet [36] 86.24 85.99 86.33 91.00
PointNet++ [37] 86.60 88.82 87.71 90.74
DGCNN [46] 87.72 89.14 88.78 91.16
3DGCN [29] 89.29 89.81 89.00 92.62
Ours 91.80 92.52 91.76 94.20
@@ 5 Conclusions
In this paper, we studied the constraints of CAD shapes and introduced its deep learning-friendly expression. Afterward, we proposed the CstNet for extracting and leveraging constraints. Additionally, we built up a multi-modal CAD shape dataset Param20K. Extensive experiment results demonstrate that our CstNet achieves state-of-the-art performance on both accuracy and robustness. Comprehensive ablation studies further validate the effectiveness of constraints, SurfaceKNN, and network backbone.
CAD shapes possess numerous attributes. In this paper, we analyze the constraints from a functional perspective. In future work, we plan to investigate additional aspects such as part design and manufacturing procedure, to achieve a deeper understanding of CAD shapes.
@@ F Overview
In Sec. G, we present the details of network architecture and data preparation in validation experiments. In Sec. H, we provide the data construction and static of Param20K dataset. Finally, Sec. I offers data process for visualization and more results of constraint acquisition and classification experiments. In Sec. J, we show more details of our constraint representation and constraint learning methods.
G Validation Experiment
G.1 Constraint Expression
@@ G.3 Model Design
H Param20K Dataset
H.1 Why build up the Param20K dataset?
The limited availability of B-Rep datasets poses barriers to deep learning on CAD shapes. Most existing CAD shape datasets consist of mesh files, such as MCB [18] and ESB [14]. While mesh files could approximate the appearance of CAD shapes, lack crucial boundary information. In contrast, B-Rep data [21, 61] serves as the native representation of CAD shapes and is therefore more suitable for dataset construction. However, labeled B-Rep datasets remain relatively scarce, for example, FabWave [1] includes only 2,133 B-Rep files. Although the ABC [20] contains a large number of B-Rep files, it remains unlabeled.
H.2 Dataset Construction
This section provide details of purification process and data acquisition of Param20K dataset.
@@ H.2 Dataset Construction
Point cloud: The point cloud data in Param20K dataset containing about 2,200 points, point constraints is also included, each point is expressed as (Coordinate, Primitive Type, Main Axis Direction, Adjacency), the meshes and point clouds are generated from B-Rep data.
H.3 Dataset Statics
This section presents the shapes count and proportion statistics for each category in the Param20K dataset. Additionally, shapes from each category are showcased.
Param20K dataset is divided into 80 % for training and 20 % for testing.